OpenStudy (lolipop_drum312):

What is − 3/8 as a decimal and as a percent? Decimal: Percent: %

OpenStudy (lolipop_drum312):

i will medal

OpenStudy (ltrout):

Well solve this equation \[\frac{ 3 }{ 8 } = \frac { x }{100}\]. Cross multiply you get 8x = 300. Divide by 8 you get 37.5. So 3/8 = 37.5 and -3/8 = -37.5. I think you can figure out the percent yeah?

OpenStudy (lolipop_drum312):

um ok

OpenStudy (ltrout):

What do you think the percent would be?

OpenStudy (lolipop_drum312):

um idk i think i divided something wrong

OpenStudy (ltrout):

You shouldn't have to divide anything to get the percent. The decimal 37.5 would be the percent aswell.

OpenStudy (lolipop_drum312):

ok so the decimal would be....would we have to use mathematics to see what that one is?

OpenStudy (ltrout):

Yes, you would have to divide 37.5 / 100 or just move the decimal 2 to the right, so .375

OpenStudy (lolipop_drum312):

ok

OpenStudy (mathmale):

What is − 3/8 as a decimal? Either on paper or on your calculator, divide -3 by 8. Do this before worrying about finding the equivalent as a percentage.

OpenStudy (mathmale):

Are you doing this math by hand or with a calculator?

OpenStudy (lolipop_drum312):

i got 0.375?

OpenStudy (lolipop_drum312):

calculator

OpenStudy (ltrout):

It's a negative fraction so your answer should end up negative by the way.

OpenStudy (ltrout):

(3/8) = 0.375, so -(3/8) should be what?

OpenStudy (lolipop_drum312):

idk sorry

OpenStudy (ltrout):

You see that negative sign infront of the second fractions ^?

OpenStudy (ltrout):

Just carry it over to the final answer of (3/8)

OpenStudy (lolipop_drum312):

um...how do we carry it over? sorry i'm terrible at math.

OpenStudy (ltrout):

Okay so you see -(3/8), don't worry about the - sign, just divide (3/8) and you get 0.375. Now you didn't do anything with the negative sign and we can't forget it, so just put it in front of 0.375 to get -0.375. Make sense?

OpenStudy (lolipop_drum312):

ooh ok thank you..so our answer is -0.375? right?

OpenStudy (lolipop_drum312):

for decimal

OpenStudy (ltrout):

Yes, for the decimal. Now just move the decimal two places to the left to get -37.5%

OpenStudy (lolipop_drum312):

ok thnk you

OpenStudy (mathmale):

"Yes, for the decimal. Now just move the decimal two places to the left to get -37.5%" Don't you mean "2 places to the right?"

OpenStudy (ltrout):

Ah yes. Meant to put right* My apologies. Thanks for spotting that @mathmale . Still early for me and tired.